Given population 15, 12, 6, 9, and 17. Suppose samples of size 3 are drawn from this population. Complete the table and draw the sketch of histogram. Describe the distribution

1
Expert's answer
2022-04-27T09:28:12-0400

1. We have population values 6,9,12,15,17, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{6+9+12+15+17}{5}=11.8"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{5}(33.64+7.84+0.04"


"+10.24+27.04)=15.76"

"\sigma=\sqrt{\sigma^2}=\sqrt{15.76}\approx3.9699"

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 6,9,12 & 27/3 \\
 \hdashline
 2 & 6,9,15 & 30/3 \\
 \hdashline
 3 & 6,9,17 & 32/3 \\
 \hdashline
 4 & 6,12,15 & 33/3 \\
 \hdashline
 5 & 6,12,17 & 35/3 \\
 \hdashline
 6 & 6,15,17 & 38/3 \\
 \hdashline
 7 & 9,12,15 & 36/3 \\
 \hdashline
 8 & 9,12,17& 38/3 \\
 \hdashline
 9 & 9,15,17 & 41/3 \\
 \hdashline
 10 & 12,15,17 & 44/3 \\
 \hdashline 
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 27/3 & 1/10 & 27/30 & 729/90 \\
 \hdashline
 30/3 & 1/10& 30/30 & 900/90 \\
 \hdashline
 32/3 & 1/10 & 32/30 & 1024/90 \\
 \hdashline
 33/3 & 1/10 & 33/30 & 1089/90 \\
 \hdashline
 35/3 & 1/10 & 35/30 & 1225/90 \\
 \hdashline
 36/3 & 1/10 & 36/30 & 1296/90 \\
 \hdashline
 38/3 & 2/10 & 76/30 & 2888/90 \\
 \hdashline
 41/3 & 1/10 & 41/30 & 1681/90 \\
 \hdashline
 44/3 & 1/10 & 44/30 & 1936/90 \\
 \hdashline
\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=11.8=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{12768}{90}-(11.8)^2=\dfrac{7.88}{3}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{7.88}{3}}\approx1.6207"